OpenStudy (anonymous):

he amount of heat needed to raise 2.0 kg of a substance by 80 K is 33 kJ. What is the specific heat of the substance? (Points : 3) 210 2,100 1,300 1.3 × 106

OpenStudy (anonymous):

Is chemestry?

OpenStudy (anonymous):

physics

OpenStudy (johnweldon1993):

\[\large Q = cm\Delta T\] Q = heat added c = specific heat m = mass \(\large \Delta T\) = change in temp rearrange and solve for C

OpenStudy (anonymous):

I dont even know where to start with this problem. I got that part right 2.0*80*33=c

OpenStudy (anonymous):

i believe its d but im confused at the same time

OpenStudy (anonymous):

no its like 33/2*80 somethinglike that right

OpenStudy (anonymous):

am i right?

OpenStudy (johnweldon1993):

\[\large C = \frac{Q}{m\Delta T}\] \[\large C = \frac{33000J}{2.0kg \times 80K}\]

OpenStudy (anonymous):

i got 206.25 for that and it wasnt one of the answer choices, I'm trying to figure out what am i missing. i know you got the 33000 by multiplying 33*1000 because thats what k stands for

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

i dont know what to do next

OpenStudy (anonymous):

please help

OpenStudy (johnweldon1993):

Well hang on a second...lets look at units lol \[\large C = \frac{Q}{m\Delta T}\] \[\large C = \frac{33000 J}{2000g \times 80K}\] We know = C = J / gk so that looks right...now \[\large C = \frac{33000}{160000gK }\] C = .20625 well that doesnt help either O.o hmm